name the property of real numbers illustrated by each equation.
2(3x-y)=6x-2y
3+(pi)=(pi)+3

Answer by jim_thompson5910(35256)

name the property of real numbers illustrated by each equation.
2(3x-y)=6x-2y
3+(pi)=(pi)+3


-------------------------------------------------------------------------------

2(3x-y)=6x-2y ... This is using the distributive property

Distributive Property

Some other examples of the distributive property


(Image Source: http://www.mathwarehouse.com/dictionary/D-words/distributive-property-diagramB.png)

Notice how the outer term is being multiplied by EVERY term inside the parenthesis and then you add up those products. A common mistake is to multiply the outer term with only the first term inside (or only a few terms inside, but not all). Be careful not to fall for this common mistake.

-------------------------------------------------------------------------------

3+(pi)=(pi)+3 ... This is using the Commutative Property of Addition

Commutative Property of Addition

This property basically states: "You can add any two numbers in any order you want"

Other examples

7+11 = 18
11+7 = 18

9+8 = 17
8+9 = 17

44+6 = 50
6+44 = 50

You can use a number line to display each example visually.

I'm sure you are quite familiar with this property at this point on an intuitive/common sense level. The "Commutative Property of Addition" is just the term mathematicians have given the idea.